Equal amounts are invested at 4%, 7%, and 8% annual interest. If the three investments yield a total of $760 annual interest, find the total investment.
Answer by ikleyn(52814)

.

Your equation is


    0.04x + 0.07x + 0.08x = 760  dollars.


Simplify 


            0.19x         = 760

              x           = 760/0.19 = 4000  dollars.


Thus each investment was  $4000;  the total investment was  $4000 + $4000 = $4000 = $12000.    ANSWER

Solved.